OpenStudy (anonymous):

What is the measure of the angle formed by the minute hand and the hour hand of a clock when the clock shows 4:35? A. 72.5° B. 90° C. 120° D. 72°

OpenStudy (anonymous):

360 ÷ 60 = 6º for each tick mark. THen, the minute hand will be at 35. The hour hand will be at 7/12 = x/5. x = 22.91 Subtract and multiply by 6. It is about A.

OpenStudy (anonymous):

is it simply 90